White light is used in a Young’s double slit experiment. Find the minimum order of the violet fringe (λ = 400 nm) which overlaps with a red fringe (λ = 700 nm).

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. Let m base th bright fringe of violet light overlaps with n base th bright fringe of red light. ∴ m * 400nm * D/d = n * 700nm * D/d ⇒ m/n = 7/4 ⇒ 7 base th bright fringe of violet light overlaps with 4 base th bright fringe of red light (minimum). Also, it can be seen that 14 base th violet fringe will overlap 8 base th red fringe. Because, m/n = 7/4 = 14/8.